MONT 107Q – Thinking About Mathematics Study/Discussion Questions on All the Pretty Horses^{1} January, 2017

1. Thinking about literary genres. A literary genre is essentially a category or style of literary composition. Genres can be defined by their content. For instance, the use of facts from science

• to project how humans will explore the currently unknown,

• to extrapolate the ways human society may develop in the future, or

• to speculate about what humans may learn from contact with alien life forms,

more or less defines the genre of science fiction. In other cases, genres are defined by their conventions and tone, and even the dimension or scope of their typical works (e.g. novels vs. short stories). (Literary critics also sometimes use the term “genre fiction” in a derogatory way to create a border around writing that is seen as more specialized, less universal and serious, with lower artistic goals, etc.) The “Western” has been an important and popular genre in American literature and film from the beginning of the 20th century to the present. Drawing on other books you have read and movies you have seen, what are the typical conventions of a “Western?” In what ways is All the Pretty Horses a “Western?” In what ways, if any, does it transcend the “Western” genre?

2. One one level, there is an obvious border between the United States and Mexico that plays an important role in this novel. But note that on the last page of All the Pretty Horses, John Grady Cole rides off into a gorgeously and poetically described sunset. What is the symbolism here? How does the whole novel relate to this symbolism? In what way does this symbolize another sort of border, or boundary?

3. “Just the facts, ma’am.” When and where does the novel take place? Give a one-page synopsis of the plot. Summarize the personality, circumstances, and outlook of each of these main characters:

• John Grady Cole–in particular how old is he? Does he seem like other people you have known at his age?

• Lacey Rawlins–same question.

• “Jimmy Blevins”–same question.

• Don Hector Rocha y Villareal.

• Alejandra.

• Duena Alfonsa.

• the police Captain in the town of Encantada (we don’t learn his full name). And by the way, what does the Spanish word encantada mean? Is that meaningful?

• the prisoner Perez.

4. In All the Pretty Horses, I think we would have to agree that the main character, John Grady Cole, grows up. He goes to Mexico with dreams about what it means to be a man and a cowboy, but what happens to those dreams as a result of the trip? How

1 Some of these questions are based on study questions from the Sparknotes page for

this novel — http://www.sparknotes.com/lit/allthepretty/study.html

does he change? What has he changed into? Would it be right to say he is a hero? What would it mean to call him a hero in the terms of the society and culture(s) portrayed in the novel?

5. Some recent literary criticism has argued that “Western” films and novels deal primarily with masculine issues and leave the concerns and experiences of women almost entirely out of the picture. Women appear only as plot devices or as sexual objects and the main focus is on the importance of masculinity, standards of male conduct, and the problem of creating an authentic male identity. To what extent do you think this is true about All the Pretty Horses? Do women have important roles? Are they different from male roles? What sorts of relationships does John Grady Cole have with women? What are his attitudes about women?

6. Before John Grady Cole leaves Texas for his journey, he sees his ex-girlfriend, Mary Catherine Barnett, and they have an awkward parting. She asks whether they can be friends but he says, “It’s just talk, Mary Catherine.” “Everything’s talk isn’t it?” she asks. “Not everything,” he answers. How does this exchange relate to conventions of masculinity, in “Westerns” or otherwise?

8. Cormac McCarthy, the author of All the Pretty Horses, makes some unusual choices about rendering conversation: he does not use quotation marks, and he presents Spanish-language utterances without translating them into English. This is certainly not an accident; it’s a conscious choice of a masterful author who is keenly attuned to the effects he is creating through his writing. What effects do these choices have on the reader and on the narrative? If some of the novel’s conversations are difficult to understand, what effect does that give? What does it mean about the novel and the world of the novel?

9. Finally, consider the title All the Pretty Horses. What does it mean? How does it relate to the plot, characters, and the novel’s important themes?